OpenStudy (anonymous):

FAN AND MEDAL If x varies inversely with y and x = 8 when y = 10, find y when x = 6. A.Y=3/20 B. y=4.8 C. y=7.5 D.y=40/3

OpenStudy (anonymous):

x = k/y 8 = k/10 k = 80 so 6 = 80/y y = 80/6 y = 40/3 ..D

OpenStudy (loser66):

yup

OpenStudy (anonymous):

thnx!!